OpenStudy (anonymous):

Using b2 − 4ac find its value given the following quadratic equation 2x2 + 7x + 3 = 0 Can someone please tell me the answer. I have tried this about 6 times and have gotten it wrong every time. Thanks in advance.

OpenStudy (anonymous):

The discriminant, b² - 4ac, is the answer. If b² - 4ac = 0, then the solution to the equation boils down to x = (-b ± 0)/2a = -b/2a, and there is only one solution to the equation, which means the equation cannot be factored by ordinary means. 2x² + 7x + 3 = 0 b² - 4ac = (7)² - 4(2)(3) = 49 - 24 = 25.

OpenStudy (anonymous):

yeaah thts the answer

OpenStudy (anonymous):

So is this the answer? b² - 4ac = (7)² - 4(2)(3) = 49 - 24 = 25 The answer can't have any commas, dollar signs, percent signs, variables, or spaces. And it says if there are extra spaces at the end that my answer will be marked wrong? I need the simplest or simplified answer

OpenStudy (anonymous):

(7)² - 4(2)(3) = 49 - 24 = 25 so this is ok for the answer or do I need the first part on there as well?

OpenStudy (anonymous):

just write it as i have it

OpenStudy (anonymous):

ok thank you so much.